OpenStudy (anonymous):

simplify: (3-2radical2)square root

OpenStudy (anonymous):

do you mean this: \[\left( 2-3\sqrt{2} \right)^{2}\]?

OpenStudy (anonymous):

\[\sqrt{3-2\sqrt{2}}\]